a plant charge q=40nc is placed in front of a circuit disc of 1m radius and at a distance of 0.01cm. the disk carries a positive charge uniformly distributed with a surface charge density of 20nc/m2. what is the force on the charge points?
1
Expert's answer
2020-05-14T09:34:06-0400

Here ,

we will find electric field"(E)" due to a charged disc at a distance of "(x=0.01\ cm=10^{-4}m)" and then

Electric force "=qE"

Given, "\sigma =20\ nC/m^2=20\times 10^{-6}C/m^2;x=10^{-4}m;" "R=1\ m" ";k=\frac{1}{4\pi \epsilon_o}=9\times 10^9" and "x<<R"

Electric field"(E)= 2\pi k\sigma(1-\frac{x}{\sqrt{x^2+R^2}})" and "x<<R" .So,

"E=2\pi k\sigma (1-\frac{x}{R})" "=2\pi \times 9\times 10^9\times 20\times 10^{-6}(1-\frac{10^{-4}}{1})\approx36\pi \times 10^4N/C"

Electric force on the charge"=qE=40\times 10^{-6}\times 36\pi \times 10^4N=45.2\ N"